# End-To-End Causal Effect Estimation from Unstructured Natural Language Data

### v1

@@ 1 Introduction
Yet, unstructured observational data presents unique opportunities for cheaper, more accessible, and potentially even better [29] effect estimation. For example, thousands of people living with diabetes choose to share their experiences with treatments on online patient forums. Some of their posts contain rich descriptions of daily lives, the drugs they have been prescribed, the treatment responses and side effects, as well as pre-treatment information like age and sex. Their posts contain their lived experiences including evidence of an outcome in an observational experiment, albeit in an unstructured form. Other potential sources of rich unstructured, observational data include newspaper classifieds, police reports, social media, and clinical reports. Despite being collected for a myriad of purposes, researchers have often turned to such data to test hypotheses since: (i) unstructured data does not require restrictive data collection designs, e.g., measurement choice, and can admit many different post-hoc analyses; (ii) the reported outcomes may reflect what matters to subjects better than standard outcome measures; (iii) value may be recouped from outcomes that would otherwise be lost; (iv) there may be more unstructured data available on underserved or marginalized populations. Figure 1 contrasts our setting with previous works using randomized or structured observational data.
@@ 1.1 Related Work
The use of natural language data in causal inference comes in different flavors: i) using text to measure confounders [21], ii) using text to measure causal effect outcomes [13], or iii) producing interpretable causal features from text [13, 5], e.g., what words are more likely to explain the cause of an event. NATURAL distinguishes itself from these lines of research in two ways: i) NATURAL does not require any curated task-specific training data (it is zero-shot), and ii) NATURAL is not interested in how the text itself, i.e., its words, relate to the causal problem —that is, we are only leveraging the model’s ability to predict the distribution of a specified variable conditional on the input text. We highlight that our work lies distinct from research at the intersection of text and causality that has studied the ability of language models to infer latent variables (that are implied but not explicitly identified in text data) [33, 12]. Rather, we require the precise specification of covariates to condition on – we view this as being crucial to creating a more direct way for an end user to verify the validity of information extracted with our approach.
@@ 3 NATURAL estimators of the ATE
Intuitively, these assumptions give NATURAL indirect access to $(X,T,Y)$ through $R$. They can be weak or strong, depending on the definition of the reports $R$. On the one hand, if reports are copies of the observational data, i.e., $R=(X,T,Y)$, then Assumption 4 is trivial to satisfy. On the other hand, if reports are all the constant, empty string, $R=\epsilon$, then Assumption 4 guarantees that we have full access to the true observational joint density function over $(X,T,Y)$, which is a strong assumption. We consider how we might satisfy these assumptions in practice in the next section. Here, we assume that they hold and develop a series of consistent estimators of the ATE.
@@ A Monte Carlo estimate over reports is given by
which further approximates $\hat{e}_{\text{N-Full}}(x)$ from the given conditional. We used eq. 7 below. We note that $\hat{\tau}_{\text{N-Full}}$ above is derived using the form of IPW, but the same estimator can be derived from OI, as shown in appendix B.
@@
$\hat{\tau}_{\text{N-OI}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\tau}_{\text{N-OI}}(X_{i},1)-\hat{\tau}_{\text{N-OI}}(X_{i},0).$ (10)
@@ 4 Implementing NATURAL estimators with Large Language Models
LLMs are trained on vast datasets of real-world data, e.g., [2], which likely contain records of data generated by processes that are consistent with Assumption 3. Because LLMs can learn well-calibrated conditionals [20], our hypothesis is that LLMs can be prompted to approximate the conditionals required by Assumption 4 for real-world causal effect questions of interest. Our LLM implementation of NATURAL estimators is built on this hypothesis to try to satisfy Assumptions 3 and 4 (Assumptions 1 and 2 must be guaranteed by a domain expert). We defer exact prompts to appendix D, a full worked example to appendix C, and a discussion of the limitations of our approach to the next section. Figure 2 summarizes our pipeline.
@@
    Filter known covariates by inclusion criteria. We are sometimes interested in ATEs over populations defined by constraints on pre-treatment covariates $X_{i}$ known as inclusion criteria. In such cases, we included a filter to enforce inclusion criteria. Managing inclusion criteria is complicated by the fact that many reports $R_{i}$ contain no or partial information about covariates that are required to verify inclusion. So, in this filtering step, our goal was to ensure that the final set of reports have non-zero probability of matching the inclusion criteria. We begin by prompting an LLM to extract the full set of covariates $X_{i}$, following constraints on the possible values each attribute can take, but we allow the LLM to extract Unknown if it is impossible for the LLM to determine the value of a covariate. We then remove reports, if any of the non-Unknown covariates are determined to fail their inclusion criteria. We found the JSON-mode made available for generation by certain LLM APIs, like GPT-4, to suffice for this task; however more involved strategies for constrained generation are also possible [43, 47].
@@
    Infer conditionals. Given $\{R_{i},X_{i}\}_{i=1}^{n}$ from the previous steps, we compute the probabilities $P_{\text{LLM}}(T=t,Y=y|R_{i},X_{i})$ by prompting an LLM that makes log-probabilities accessible, like LLAMA2. Specifically, we ask an LLM to answer questions about $T,Y$ given access to $R_{i},X_{i}$, and we score every possible answer $T=t,Y=y$ using the LLM log-probabilities. We exponentiate and renormalize these scores across the space of possible realizations to obtain a valid probability distribution.
To manage inclusion criteria, our LLM pipeline requires three additional technical conditions, which we discuss in more detail in appendix G: (i) inclusion criteria define a box, i.e., a separate criterion is specified for each covariate dimension, (ii) the event that the set of Unknown in (iv) satisfy their inclusion criteria is conditionally independent of the report and the known covariates given the event that the knowns satisfy their criteria, and (iii) the set of known and unknown covariates is conditionally independent of the treatment and outcome given the report and the covariate values. We used these to avoid an expensive estimation or a data-wasteful rejection sampling step, but these assumptions are not required in principle by NATURAL estimators.
@@ 5 Limitations and Broader Impact
  • Selection Bias. Results might be biased towards individuals’ choice of reporting an outcome given their experience with the treatment. Luckily, outcome missingness is a widely studied problem in causality research, see e.g., how to test [8] or how to mitigate [28] it. Note, however, that solutions will often accumulate assumptions on top of NATURAL and should always be critically evaluated by practitioners. Finally, apart from individuals’ choice of reporting, selection bias might also arise from which individuals participate in online forums, i.e., our framework is only capable of estimating local ATEs —external validity is not guaranteed a priori.
@@ 6 Empirical Evaluation
Ground Truth $\mathbf{10.11}$ [NCT03987919, 16] $\mathbf{-14.7}$ [NCT03191396, 7] $\mathbf{28.3}$ [NCT03828539, 34] $\mathbf{41.00}$ [NCT02191579, 36]
@@
We present our estimated ATE and its RMSE on the synthetic datasets in table 1. Further, we evaluate two trained baselines, which use a Bag-of-Words model and a sentence encoder respectively, to train representations of text data with their labels. Here, for each attribute in the set of covariates, treatments, and outcomes, we train a MLP model with 5-fold cross validation to predict that attribute. We then use these predicted attributes as a tabular dataset of samples that can be plugged into any causal inference estimator. We find that our methods are competitive with or outperform these baselines, despite not being trained with any labels. In particular, the sentence encoder baseline collapsed to an ATE of zero for Hillstrom, having learned the constant predictor for outcomes.
@@ 7 Conclusion
In this work, we introduced NATURAL, a family of text-conditioned estimators, to automate treatment effect estimation using free-form text data. We demonstrated NATURAL’s efficacy with six synthetic and real datasets for systematic evaluation of its pipeline. We exposed the ability of LLMs to extract meaningful conditional distributions over structured variables and how their combination with classical causal estimators can predict real-world causal effects with remarkable accuracy. Given this promising performance, exciting directions for future work include (i) incorporating automatic prompt tuning methods into the pipeline, (ii) extending our methods to real-valued treatments, outcomes and covariates, (iii) exploring whether our assumptions can be weakened, (iv) exploring other domains in applied research, e.g., social sciences, (v) performing a more extensive evaluation of NATURAL on different study designs to better understand what type of treatments, outcomes, and reports show better or worse practical performance with NATURAL or (vi) deploying the pipeline to test hypotheses at even larger scales.
@@ Appendix G Inclusion Criteria conditioned Estimator
For example, if $\tau(I)$ can estimated by the IPW estimator,
@@
because the $P(T=1|X=x,X\in I)=P(T=1|X=x)$. Throughout this section, we operate under Assumptions 3 and 4 and assume that the LLM gives us access to the true data-generating conditionals.
@@
It is possible to directly estimate eq. 23 with Monte Carlo samples from $P(R)$ and an approximation of $P(X\in I|R)/P(X\in I)$ using an LLM, but the latter is computationally very expensive. Alternatively, one can use rejection sampling to simulate $R|X\in I$ by sampling $X_{i}\sim P(X|R_{i})$ and rejecting $R_{i}$ if $X_{i}\notin I$. This step could be very wasteful for precious data. In fact, in our experiments with the Semaglutide vs. Tirzepatide dataset, we found that using GPT-4 Turbo for sampling $X_{i}\sim P(X|R_{i})$ resulted in a very peaky distribution with little diversity in the extractions, even after increasing its temperature argument as much as possible without sacrificing generation quality. As a result, after rejecting $R_{i}$ if $X_{i}\notin I$, we were left with a very small number of reports, that would be infeasible to plug into any estimator. Therefore, we devise a few assumptions that allow us to estimate eq. 21 without these approaches.

### v2

@@ 1 Introduction
Yet, unstructured observational data presents unique opportunities for cheaper, more accessible, and potentially even better [32] effect estimation. For example, thousands of people living with diabetes choose to share their experiences with treatments on online patient forums. Some of their posts contain rich descriptions of daily lives, the drugs they have been prescribed, the treatment responses and side effects, as well as pre-treatment information like age and sex. Their posts contain their lived experiences including evidence of an outcome in an observational experiment, albeit in an unstructured form. Other potential sources of rich unstructured, observational data include newspaper classifieds, police reports, social media, and clinical reports. Despite being collected for a myriad of purposes, researchers have often turned to such data to test hypotheses since: (i) unstructured data does not require restrictive data collection designs, e.g., measurement choice, and can admit many different post-hoc analyses; (ii) the reported outcomes may reflect what matters to subjects better than standard outcome measures; (iii) value may be recovered from outcomes that would otherwise be lost; (iv) there may be more unstructured data available on underserved or marginalized populations. Figure 1 contrasts our setting with previous works using randomized or structured observational data.
@@ 1.1 Related Work
The use of natural language data in causal inference comes in different flavors: i) using text to measure confounders [23], ii) using text to measure causal effect outcomes [15], or iii) producing interpretable causal features from text [15, 6], e.g., what words are more likely to explain the cause of an event. NATURAL distinguishes itself from these lines of research in two ways: i) NATURAL does not require any curated task-specific training data (it is zero-shot), and ii) NATURAL is not interested in how the text itself, i.e., its words, relate to the causal problem —that is, we are only leveraging the model’s ability to predict the distribution of a specified variable conditional on the input text. We highlight that our work lies distinct from research at the intersection of text and causality that combines text and numerical or tabular data [14], the latter of which may be unavailable or incomplete in settings involving neglected diseases, unrecorded abortions, or illicit drug use. Other work has also studied topic modelling approaches [2] or the ability of language models to infer latent variables (that are implied but not explicitly identified in text data) [36, 13]. Rather, we require the precise specification of covariates to condition on – we view this as being crucial to creating a more direct way for an end user to verify the validity of information extracted with our approach.
@@ 3 NATURAL estimators of the ATE
Intuitively, these assumptions give NATURAL indirect access to $(X,T,Y)$ through $R$. They can be weak or strong, depending on the definition of the reports $R$. On the one hand, if reports are copies of the observational data, i.e., $R=(X,T,Y)$, then Assumption 4 is trivial to satisfy. On the other hand, if reports are all the constant, empty string, $R=\epsilon$, then Assumption 4 guarantees that we have full access to the true observational joint density function over $(X,T,Y)$, which is a strong assumption. In other words, it requires a way to simulate trial outcomes unconditionally (without any data). We consider how we might satisfy these assumptions in practice in the next section. Here, we assume that they hold and develop a series of consistent estimators of the ATE.
@@ A Monte Carlo estimate over reports is given by
which further approximates $\hat{e}_{\text{N-Full}}(x)$ from the given conditional. We used eq. 7 below. We note that $\hat{\tau}_{\text{N-Full}}$ above is derived from IPW, but can also be derived from OI, as shown in appendix B.
@@
$\hat{\tau}_{\text{N-OI}}=\frac{1}{n}\sum_{i=1}^{n}\hat{\tau}_{\text{N-OI}}(X_{i},1)-\hat{\tau}_{\text{N-OI}}(X_{i},0)$ (10)
@@
Inclusion Criteria conditioned ATE. We are sometimes interested in ATEs over populations defined by constraints on pre-treatment covariates $X_{i}$, known as inclusion criteria and denoted by $I$. This conditional ATE is $\tau(I)=\mathbb{E}[Y(1)-Y(0)\mid X\in I]$ and satisfies the following identity.
$\tau(I)$ $=\mathbb{E}_{R}\left[\frac{P(X\in I|R)}{P(X\in I)}\mathbb{E}_{X,T,Y}\left[\frac{TY}{e(X)}-\frac{(1-T)Y}{1-e(X)}\middle|X\in I,R\right]\right].$ (11)
The conditional version of NATURAL IPW can be estimated by filtering out reports where $P(X\in I|R_{i})=0$, sampling $X_{ij}\sim P(X|R_{i},X\in I)$ i.i.d., and finally weighting datapoints by the relative likelihood of matching the inclusion criteria given the report:
$\hat{\tau}(I)=\sum_{i=1}^{n}\left[\frac{P(X\in I|R_{i})}{\sum_{i=1}^{n}P(X\in I|R_{i})}\sum_{j=1}^{m}\frac{1}{m}\mathbb{E}_{T,Y\mid X_{ij},R_{i}}\left[\frac{TY}{e(X_{ij})}-\frac{(1-T)Y}{1-e(X_{ij})}\right]\right],$ (12)
where the inner expectation is estimated similar to eq. 6. A complete derivation for eqs. 11 and 12 and related discussion are included in appendix G. In practice, we took $m=1$.
@@ 4 Implementing NATURAL estimators with Large Language Models
LLMs are trained on vast datasets of real-world data, e.g., [3], which likely contain records of data generated by processes that are consistent with Assumption 3. Because LLMs can learn well-calibrated conditionals [22], our hypothesis is that LLMs can be prompted to approximate the conditionals required by Assumption 4 for real-world causal effect questions of interest. Our LLM implementation of NATURAL estimators is built on this hypothesis to try to satisfy Assumptions 3 and 4 (Assumptions 1 and 2 must be guaranteed by a domain expert). We defer exact prompts for LLM inference to appendix D, a full worked example to appendix C, and a discussion of the limitations of our approach to the next section. Figure 2 summarizes our pipeline.
@@
    Filter known covariates by inclusion criteria. For ATEs conditioned on inclusion criteria, as in our real-world datasets, we included a filter to enforce these criteria. Managing inclusion criteria is complicated by the fact that many reports $R_{i}$ contain no or partial information about covariates that are required to verify inclusion. So, in this filtering step, our goal was to ensure that the final set of reports have non-zero probability of matching the inclusion criteria. We begin by prompting an LLM to extract the full set of covariates $X_{i}$, following constraints on the possible values each attribute can take, but we allow the LLM to extract Unknown if it is impossible for the LLM to determine the value of a covariate. We then remove reports, if any of the non-Unknown covariates are determined to fail their inclusion criteria. We found the JSON-mode made available for generation by certain LLM APIs, to suffice for this task; however more involved strategies for constrained generation are also possible [46, 50].
@@
    Infer conditionals. Given $\{R_{i},X_{i}\}_{i=1}^{n}$ from the previous steps, we compute the probabilities $P_{\text{LLM}}(T=t,Y=y|R_{i},X_{i})$ by prompting an LLM that makes log-probabilities accessible. Specifically, we ask an LLM to answer questions about $T,Y$ given access to $R_{i},X_{i}$, and we score every possible answer $T=t,Y=y$ using the LLM log-probabilities. We exponentiate and renormalize these scores across the space of possible realizations to obtain a valid probability distribution.
  3. (vii)
    Weight reports according to inclusion criteria match. Similar to item (vi), we compute $P_{\text{LLM}}(X_{i}\in I|R_{i})$ to obtain the weights in eq. 12, by prompting an LLM with descriptions of the inclusion criteria that must be satisfied and each report $R_{i}$. It may be possible to skip this weighting step under additional structural assumptions on the data. These assumptions as well as experimental results without the weighting are included in appendix G.
@@ 5 Limitations and Broader Impact
  • Selection Bias. Results might be biased towards individuals’ choice of reporting an outcome given their experience with the treatment. Luckily, outcome missingness is a widely studied problem in causality research, see e.g., how to test [9] or how to mitigate [31] it. Note, however, that solutions will often accumulate assumptions on top of NATURAL and should always be critically evaluated by practitioners. Finally, apart from individuals’ choice of reporting, selection bias might also arise from which individuals participate in online forums, i.e., our framework is only capable of estimating local ATEs —external validity is not guaranteed a priori. We demonstrate this challenge in section 6 by simulating systematic bias in synthetic settings.
@@ 6 Empirical Evaluation
N-MC OI $5.89\pm 1.03$ $4.28$ $-8.23\pm 0.94$ $6.54$ $25.62\pm 0.51$ $2.72$ $46.20\pm 1.94$ $5.55$
N-MC IPW $5.62\pm 0.85$ $4.81$ $-7.10\pm 0.94$ $7.66$ $26.65\pm 1.44$ $3.19$ $46.52\pm 1.92$ $5.85$
N-OI $4.84\pm 1.19$ $5.39$ $\mathbf{-16.57\pm 1.06}$ $\mathbf{2.15}$ $\mathbf{29.05\pm 1.77}$ $\mathbf{1.92}$ $44.67\pm 1.56$ $3.99$
N-IPW $\mathbf{9.06\pm 0.69}$ $\mathbf{1.26}$ $-12.54\pm 0.86$ $2.33$ $25.64\pm 0.40$ $2.68$ $\mathbf{42.53\pm 2.07}$ $\mathbf{2.57}$
Ground Truth $\mathbf{10.11}$ [NCT03987919, 18] $\mathbf{-14.7}$ [NCT03191396, 8] $\mathbf{28.3}$ [NCT03828539, 37] $\mathbf{41.00}$ [NCT02191579, 39]
@@
We present our estimated ATE and its RMSE on the synthetic datasets in table 1. Further, we evaluate two trained baselines, which use a Bag-of-Words model and a sentence encoder respectively, to train representations of text data with their labels. Here, for each attribute in the set of covariates, treatments, and outcomes, we train a MLP model with 5-fold cross validation to predict that attribute. We then use these predicted attributes as a tabular dataset of samples that can be plugged into any causal inference estimator. We find that our methods are competitive with or outperform these baselines, despite not being trained with any labels. In particular, the sentence encoder baseline collapsed to an ATE of zero for Hillstrom, having learned the constant predictor for outcomes. IPW (Structured) is an oracle estimator, which assumes full access to structured data. Selection-biased N-IPW demonstrates the challenge of ATE estimation in the presence of bias which was systematically simulated as a function of covariates “channel" and “zip code" for Hillstrom and “age" for Retail Hero.
@@ 7 Conclusion
In this work, we introduced NATURAL, a family of text-conditioned estimators, to automate treatment effect estimation using free-form text data. We demonstrated NATURAL’s efficacy with six synthetic and real datasets for systematic evaluation of its pipeline. We exposed the ability of LLMs to extract meaningful conditional distributions over structured variables and how their combination with classical causal estimators can predict real-world causal effects with remarkable accuracy. Given this promising performance, exciting directions for future work include (i) incorporating automatic prompt tuning methods into the pipeline, (ii) extending our methods to real-valued $(X,T,Y)$, (iii) exploring whether our assumptions can be weakened, (iv) exploring other domains in applied research, e.g., social sciences, (v) performing a more extensive evaluation of NATURAL on different study designs to better understand what type of treatments, outcomes, and reports show better or worse practical performance with NATURAL or (vi) deploying the pipeline to test hypotheses at even larger scales.
@@ Appendix B Deriving NATURAL Full from Outcome Imputation estimator
    Weight reports according to inclusion criteria match. We also used LLAMA2-70B to compute the weighting terms described in item (vii) of section 4. Concretely, we constructed a prompt describing the inclusion criteria listed in item 4 of the experiment design above, followed by a report, $R_{i}$ and an instruction asking the LLM to determine whether all the described criteria are met. We then scored the possible answers, ["No", "Yes"], exponentiated and renormalized them to obtain $P(X\in I\mid R_{i})$. We marginalized over reports to compute the denominator, $P(X\in I)$, in the weight. The contribution of each report to the ATE estimates was weighted by this relative likelihood of matching the in inclusion criteria of the experiment given the report.
  8. 8.
@@ F.2 Balancing property of propensity scores
F.3 Sensitivity Analysis
While we rely on domain expertise to define the confounder set for each setting such that necessary causal assumptions are satisfied, unobserved confoundedness remains a challenge. We followed strategies in Lu and Ding, [29] to analyze the sensitivity of our ATE estimates to the degree of unobserved confoundedness. Specifically, the idea is to introduce sensitivity parameters,
$\epsilon_{0}(X)=\frac{\mathbb{E}[Y(0)|T=1,X]}{\mathbb{E}[Y(0)|T=0,X]}\text{ and }\epsilon_{1}(X)=\frac{\mathbb{E}[Y(1)|T=1,X]}{\mathbb{E}[Y(1)|T=0,X]}$
Table 5: NATURAL IPW for Semaglutide vs. Tirzepatide is robust to the degree of unobserved confoundedness shown below.
$\varepsilon_{0}$ $\varepsilon_{1}$ 1.00 1.05 1.10 1.15 1.20 1.25
1.00 9.45 6.51 3.85 1.41 -0.83 -2.88
1.05 8.51 5.58 2.91 0.47 -1.76 -3.82
1.10 7.58 4.64 1.97 -0.46 -2.70 -4.75
1.15 6.64 3.71 1.04 -1.40 -3.63 -5.69
1.20 5.71 2.77 0.10 -2.33 -4.57 -6.62
1.25 4.77 1.84 -0.83 -3.27 -5.50 -7.56
which quantify the degree of unobserved confoundedness. In our case, sensitivity parameters are the density ratio between likelihood of each potential outcome in the treated vs. untreated group. The ATE estimate is non-increasing in the sensitivity parameters. So, for positive ATEs, we are looking for the largest sensitivity parameters that maintain the positivity of the ATE, which tells us the degree of unobserved confoundedness that an estimator is robust to. Table 5 shows that the direction (sign) of NATURAL IPW ATE estimates for the Semaglutide vs. Tirzepatide dataset change from positive to negative at large values of sensitivity parameters, implying that they are robust to large degrees of unobserved confoundedness.
We further investigated the importance of each confounder as suggested by the leave-one-covariate-out approach in Section 4 of Lu and Ding, [29], by dropping that covariate as if it is an unobserved confounder and measuring the corresponding worst-case (over all possible values of the remaining covariates) sensitivity parameters with the remaining covariates. Figure 11 shows contour lines depicting NATURAL IPW ATE estimates at different values of sensitivity parameters for all our real-world datasets. Orange regions denote sensitivity parameter values for which the sign of the estimate ATE does not change, while blue regions denote the values that flip the direction of the estimate. Stars show sensitivity parameters for each covariate in the different settings. Hence, this estimator is sensitive to the covariate set and all covariates are important in the case of Semaglutide vs. Tirzepatide, while most are important for Erenumab vs. Topiramate. The direction of ATE estimates is not sensitive to covariates in the case of Semaglutide vs. Liraglutide and OnabotulinumtoxinA vs. Topiramate, implying that the estimated causal effects could only be explained away by an unobserved confounder that is stronger than all observed confounders.
(a) Semaglutide vs. Tirzepatide
(b) Semaglutide vs. Liraglutide
(c) Erenumab vs. Topiramate
(d) OnabotulinumtoxinA vs. Topiramate
Figure 11: NATURAL IPW ATE estimates for different values of sensitivity parameters in each real setting. The orange (blue) region corresponds to ATEs with the same (flipped) sign as our estimates.
@@ Appendix G Inclusion Criteria conditioned Estimator
For example, $\tau(I)$ can estimated by the IPW estimator,
@@
because the $P(T=1|X=x,X\in I)=P(T=1|X=x)$ for all $x\in I$. Throughout this section, we operate under Assumptions 3 and 4 and assume that the LLM gives us access to the true data-generating conditionals.
@@
We prompted LLAMA2-70B with descriptions of the inclusion criteria and each report to estimate $\frac{P(X_{i}\in I|R_{i})}{P(X_{i}\in I)}$, similar to other conditional distributions described in section 4, and marginalized over reports for the denominator.
It is also possible to avoid this weight computation at the expense of additional structural assumptions on the data. We discuss these conditions below and show corresponding results for NATURAL estimators in table 6.
@@
Tuned Held-out
[…]
Ground Truth $\mathbf{10.11}$ [NCT03987919, 18] $\mathbf{-14.7}$ [NCT03191396, 8] $\mathbf{28.3}$ [NCT03828539, 37] $\mathbf{41.00}$ [NCT02191579, 39]
Table 6: ATE estimates on real datapoints that are filtered according to inclusion criteria but not weighted by the relative likelihood that they meet the inclusion criteria of the experiment given the report. Best performing NATURAL estimators fall within $3$ percentage points of their corresponding ground truth clinical trial ATEs. Possible ATE values lie between $-100$ and $100$.